OpenStudy (anonymous):

Find a rational number and an irrational number that are between 7.7 and 7.9. Include the decimal approximation of the irrational number to the nearest hundredth.

OpenStudy (anonymous):

@iGreen can you help me please?

OpenStudy (anonymous):

@Gabylovesyou

OpenStudy (igreen):

Rational numbers can be expressed as fractions.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so how would i express this very one?

OpenStudy (igreen):

\(7 \dfrac{8}{10}\) is one. It simplifies to 7.8

OpenStudy (anonymous):

ok so then 7.8 would be the irrational number in between 7.7 and 7.9 right

OpenStudy (igreen):

Rational, not Irrational.

OpenStudy (anonymous):

oh ok

OpenStudy (igreen):

7.797979797979 is irrational.

OpenStudy (igreen):

It can't be expressed as a fraction because it has repeating decimals. It would actually be 7.79 with a line on top of 79.

OpenStudy (anonymous):

ok that make sense

OpenStudy (anonymous):

Reina and Sam are multiplying (84)5 (73)9. Reina's Work Sam's Work (84)5 (73)9 = 84 + 573 + 9 = 89712 (84)5 (73)9 = 84⋅573⋅9 = 820727 Is either of them correct? Explain your reasoning.